Tech Staff Meeting Notes July 31, 2006

Current Top Priorities are:NMAD Node BB upgrade, Station Maintenance, and ET Stations. (also would like to fit GNAR rebuild in there somewhere).

Annual Fish Fry is Friday, August 25.

There are four levels of priority.
PrioritiesWish to try to get these done (or make progress on) before the next meeting (yeah right)
Second TierFit this in too if an opportunity arises. May be good to do instead of priorities but check with Mitch. Likely to be a priority in a previous and/or subsequent month
As Time PermitsFit these in when you can but not instead of first or second priorities
OtherStuff we hope to get to and don't want to forget about

This redistributed task list agreed to at the November meeting. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything.

Greg's August ToDo List is incorporated into this page.


BB network upgrade.
Huddle tests are in progress and several sensors were blessed by Greg. See Mitch's PDF webpage for results.
GNAR will need to include moving our equipment from the classroom to a different room.
HICK was upgraded but has gps problems and a large dc offset.
PENM and HENM will work with 20' mast. On deck for upgrade. Have tested sensors.
Jim is getting unused freewaves from Wurth.

These are notes from James with some mods by Mitch to update based on this meeting:

MACC/CJTR - These are our two most distant sites (N Little Rock, AR, Park Hills, MO). We normally only get to each once a year, and they'll both need standard maintenance (batteries, case inspection, etc). I'll need Chris McGoldrick along for CJTR since he knows the layout there. Bill can help with MACC. Both these projects should likely wait until July.
MAIR - Our steel enclosure at the Mississippi County Airport is rusting and needs to be replaced. This will be a two man job, at least. Bob needs to order the replacement enclosure, first, so we'll need to let you know when we're ready to proceed. This should be done before August. We've also got a lead on wireless Internet based in East Prairie, MO, and I've called to inquire about it. I'm still waiting for an email with equipment details before determining if we can get similar equipment for cheaper. Bob will need to give approval before we pursue this as a means of getting MAIR online.
RLAP - Our receiver at the Reelfoot Air Park needs to be placed back inside the terminal building. This project can wait until later this year, but should be completed before November. Also needs new air cells.
CERI - We need to get an antenna setup here at CERI for GPS research and testing. Looks like roof of patio or fdl is best choice. James and Greg are investigating.

John Paul has a project to install two new near field stations at reelfoot.


The map shows stations with yellow triangles being new ones.
Mt Gibbs Node is complete. Al is working on getting the remote reset working through the firewall.
May need to install additional filters at the receive end to reduce getting interference or intermod or whatever is clobbering the data occassionally.
SMNC needs to have pins verified and adsend.d verified; gain channel is on the vertical.
SMNC4done (see note above)
BHT8move (eventually)
Can we figure out a way to give the state park people some kind of wiggle display (waveviewer is not a good option and they would like something closer to real-time than heli_ewII).
ETT is repaired. Has some breaks in the coax. Should probably replace entire length.
Flat gap, Van hill, and LRVA are on deck.
Need vehicle registration and insurance documents. Either the originals or something that says they are on file at State. Not a big deal in TN but sure is in other states. John as this odd aversion to jail cells. Chris is investigating.
TZTN needs a gate and have episensors swapped. The one there now belongs to SLU.
TRYN keeps blowing a 1 amp fuse.

Bill is getting familiarized with these so he can take them over. Continued problems with mss100's losing connectivity (SMOC: simple matter of connectivity). The good news is Jim verified that the stations are still working when SMOC is lost.
John upgraded gps firmware at 2 of 3 gps. Mitch doesn't recall which ones. John says Knoxville and 2 in charleston still need upgraded. This is the list from before:
SWET-3H52, ADSC-5657, and CSTN-5637 need to have the leap second corrected. SHTN-5648 and TRSC-5662 might also but need a site visit to verify.
Mitch needs to contact UKY tech to get a site visit to MVKY.
ATTN is having firewall problems.
MSS power cycle at ADSC didn't work. John will need to visit.
Steve thinks John needs a new laptop. John doesn't think he does.
CVTN and ASAR optical isolation. Need to upgrade to year 3 hardware. Hardware ready. James brought back some nice pictures from CVTN to help figure out how to do it (e.g. the conduit is way too small).

Routine Maintenance.

Unscheduled Maintenance
Jim and Bill fixed a bunch of these. Mitch isn't sure which. Here's Holly's list from the last meeting:
LNXT Node:
NMAD Node:
MKTA Node:

SWET Upgrade DONE.

Second Tier Priorities

Bill is building some cabinets for Smalley to go to Argentina.

Need to have the pickup serviced and the Suburban rear end looked at. Done.

Greg is installing and configuring power for new Dell rack mounts. Done.

Need to follow up on ATV training.

NCCU Gordana Vlahovic is getting a cmg6td to install at North Carolina Central University. Invoice paid. Still owe here a vault (John has it) and installation advice. Need to get a public display. Scream to Scream from the datalogger pc to the display pc would be best. Don't forget to NOT allow remote config on display scream.

Steiner Brain Dump Seminar Gregs list is starting to take shape. While anyone is invited its geared to tech staff. Schedule will vary and be weather dependent to avoid field work interruptions. Will start within a few weeks.
How do we record this for posterity? Video would be enormous in size (potential of 104 hours or more of video). Audio might be useful but would not make a very good reference. Mitch thinks written is best preferably in html or pdf. Perhaps have someone (who) take everyones notes and combine them with Greg's notes and make a web page. Audio/video cache to help clarify might be useful for whoever does this.

FDL. Mitch needs to yank Paul's chain to get an account number for new laptops.

As Time Permits

Mitch needs to write an ew module to monitor the environmental channels newly installed at the nodes. Done. Its a shell script that runs twice a day. Computes several minute averages. Emails when there's a problem. Its is /gaia/home/mwithers/Projects/ENVMonitoring.

Need to come up with a good way for John to view real data (not helicorders) on is pc over dsl. Mitch thinks he did this but can't remember what the solution was.

Generators Bill is finishing up the enclosure. Need to clean off the coils.
Done and it works well. Need to keep an eye on the weeds that still manage to sneak through the ground cover.

Inventory Sid has Ketchum's INV program installed on smeagol. Jim got initial training. Need to show Bill how to use it. Jim will be the lead and Bill will do initial update. Routine updates can be done by whoever does the work (requires usernames and passwords in mysql). Not sure yet how best to handle dataless seed production. Mitch needs to learn the admin tasks. Startup script is /gaia/smeagol/local/bin/runINV.


Many of the air cells in storage are ancient and bad. Need to go through them and dispose of the old. Greg has method for safe and legal disposal.

I-40 Bridge.
Boreholes are drilled and more or less secured. Need to do prep for new instrumention at HDBT and take care of drainage now that a big pile of stone and dirt was piled around the station. Greg's theory is lightning is clobbering gps. Has suppression constructed and ready to install.
Need a better pc at AZTN now that we have equivalent of two nodes plus going through it.

Got the modem and internet works. Need to integrate into CERI central processing.

MPHB. Withers needs to determine current orientation so we can pull it and orient correctly. Helicorders look strange. Need to verify the data.

VBMS. Vault is done. Contractor work is complete. Alena says sensors might be available in a month but NEIC staff is currently tied up with carribean net (tough job but someone has to do it).

Langston Zack. Greg's list said to test accelerometers and mass produce hardware. This is wayback burner. Mitch will check with Chuck to verify that its even going to be done anymore. Mitch asked and its unknown. Strike from list.